Widget Co. was founded at the start of 2016 by issuing 10M shares at $187.50 per share and $1,200M of long term debt. All data in the table, except actual price per share, is in millions. No assets were sold during the first two years of operation. Remember when and where to include marketable securities and the current portion of long term debt in your intermediate calculations. Use this information and the table below to answer problems 28-30 Net Income Dividends Cash Marketable Sec. Acct. Rec. Inventory Gross PPE ccumulated Dep. Net PPE Acct. Payable CP LT Debt LT Debt 2016 250 125 100 100 300 300 3000 200 2800 400 2017 300 150 120 120 300 360 3600 500 3100 400 100 1400 200 100 Shares Out $ 20.00$ 22.00 28. Between December 2016 and December 2017, Widget Companys market capitalization: a. Increased by $332M b. Increased by $200M C. Increased by $156M d. Decreased by $156M e. Decreased by $200M 29. At the end of 2016, Widget Companys P/E is closest to: a. 7.2 b. 8.0 c. 14.4 d. 15.0 e. 16.0 30. During 2017, Widget Company HINT: BOY SE NI DIVS Equity Issued Equity Repurchased EOY SE 23 a. Net borrowing was $200M and shares outstanding were repurchased at a price of $24 per share b. Net borrowing was $200M and the company issued shares at a price of $25 per share c. Net borrowing was $300M and shares outstanding were repurchased at a price of $24 per share d. Net borrowing was $300M and shares outstanding were repurchased at a price of $25 per share e. Net borrowing was $300M and the company issued shares at a price of $24 per share
Answer #1

28.Market Capitalisation in 2016 = Number of Shares*Price per Share

= 100*20 = $2,000M

In 2017 = 98*22 = $2,156M

Increase in Market Capitalisation = 2,156 – 2,000 = 156 M

Hence, the answer is c.

29. P/E ratio = price per share/earnings per share

Earnings per share = Net Income/number of shares

= 250/100 = $2.5

P/E Ratio = 20/2.5 = 8

Hence, the answer is b.

30.Borrowings = Debt raised = 1400+100 – 1200

= $300M

Calculation of Equity Funds at the beginning:

Cash + Marketable securities + acct. Rec. + Inventory + Net PPE – Acc. Payables – LT Debt

= 2,000M

At the end:

= 2,100 M

Beg Equity Funds + Net Income – Dividends = Equity Repurchased – Ending Equity Funds

2,000 + 300 – 150- Equity Repurchased = 2,100

Equity Repurchased = 50M

Number of shares = 2M

Price per share = $25

Hence, the correct answer is d.